The sum of the present ages of Eric and his father is 58 years. In 10 years, his father will be twice as old as Eric will be at that time. Find their present ages.
Answer by ikleyn(53762)

.

    E + F = 58             (1)   (the sum of current ages)

    F + 10 = 2*(E + 10)    (2)    (in 10 years)


From equation (1), express F = 58 - E  and substitute it into equation (2), replacing F there.


    (58 - E) + 10 = 2E + 20

     58 + 10 - 20 = 2E + e

     48           = 3E

     E            = 48/3 = 16.


ANSWER.  Eric is 16 years old.  The father is 58-16 = 42 years old.

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.